### v1

# Latch, Spring and Release: The Efficiency of Power-Amplified Jumping

Abstract
Many small animals, particularly insects, use power-amplification to generate rapid motions such as jumping, which would otherwise not be possible with the standard power density of muscle. A common framework for understanding this power amplification is Latch-Mediated, Spring Actuated (or LaMSA) jumping, in which a spring is slowly compressed, latched in its compressed state and the latch released to allow jumping. Motivated by the jumps of certain insect larvae, we consider a variant of this in which the latching occurs externally via adhesion to a substrate that is quickly released for jumping. We show that the rate at which this adhesion is lost is crucial in determining the efficiency of jumping and, indeed, whether jumping occurs at all. As well as showing how release rate should be chosen to facilitate optimal jumping, our analysis suggests that it is possible to control jumps even once the latch has been set.
preprint: APS/123-QED
Power amplification is an important strategy for generating fast motion in both the engineering and natural sciences: an archer stores energy slowly in the string of a bow that allows the arrow to fly faster when released than would have been the case if it had simply been thrown [1]. Similarly, plants such as the Venus flytrap [2] and insects such as the family of ‘click’ beetles [3] are able to generate rapid motions or jumps by storing energy slowly in an effective spring that is released quickly when required.
A key ingredient for exploiting this power amplification is being able to control the release. An internal ‘latch’ is the common way of understanding this with a latching mechanism being explicitly identified in the click beetles, for example [3]. The integration of a motor, spring and latch to generate rapid motion is a generic strategy that has been categorized under the umbrella of ‘Latch-Mediated, Spring Actuated’ (or LaMSA) mechanisms [4]. The spring is typically loaded at low power by the motor and then latched, ready to be released when desired (for example, when an insect needs to avoid a predator). While this mechanism is very effective at releasing energy quickly when required, as typically presented, it suggests very little control over the way in which the resulting motion occurs — once the spring is loaded, how can one control features of the motion such as the jump height (for a click beetle) or the speed of the arrow (for an archer)?
Latch mechanisms are usually embodied in the elastic body itself as for the click beetles [3], but also in artificial jumping systems [5, 6]. However, in other cases another body (most often a substrate) serves as an anchor and the effective latch is the connection between the jumper and substrate. This unusual ‘external latch’ behavior is exhibited by some larval beetles [7]. It is understood that the larva attaches its legs to the substrate and then arches its body dorsally by muscle contraction. Because of the attachment to the substrate, the arching of the larva’s back does not significantly change the shape but effectively stores elastic energy in the elastic components of its body (the spring) while its legs grip the substrate (thereby forming the latch). When the legs begin to lose contact with the substrate, the constraint imposed by the latch is released and the body is able to return to its (now curved) natural state. In the process, the stored elastic energy is transferred to kinetic energy of jumping. The combination of relaxation to a curved shape and jumping is shown in fig. 1a.
Figure 1: Release rate control of externally-latched jumping. (a) Jumping behavior in larval beetles (Laemophloeus biguttatus). Adapted from Bertone et al. [7] (equally spaced frames show instants between $t=0,100\penalty 10000\ \text{ms}$ after the initation of jumping.) (b-c) A curved elastic strip is initially indented at its center by a pressing finger (b) and a hanging weight (c). A low release rate does not result in jumping ((b), equally timed frames between up to $867\penalty 10000\ \text{ms}$ after first motion), while quick release does ((c) with equally timed frames up to $64\penalty 10000\ \text{ms}$ after first motion). Unlatching times in the time composite images (a-c) are $t_{u}=(5.5,683,1)\penalty 10000\ \text{ms}$. Inset in panel (b): Schematic representation of the shell in the $xz$-plane at three different configurations; in absence of external forces (other than its own weight) in dashed blue, when it is loaded at the center (in solid blue), and when the loading is partially released (orange).
To better understand this jumping mechanism, we consider confining a naturally curved elastic strip by poking it down at its center (mimicking the legs’ grip on the substrate) and then releasing it. Figure 1b shows a superposition of the states that the arch goes through when the confining force is removed relatively slowly — the strip goes from being approximately flat to being naturally curved, but does not appear to lose contact with the substrate. However, if the confining force is removed more quickly — for example by hanging a weight below the arch via a thread that can readily be cut (as in fig. 1c) — then the strip is observed to jump from the substrate. Clearly, the strip stores elastic potential energy when it is indented that is then converted to kinetic energy. More surprisingly perhaps, the efficiency of this conversion depends on how quickly the indenting force is released. This raises the question of how the speed of release changes the efficiency of the conversion from elastic potential energy to kinetic energy of jumping.
The jumping larvae studied by Bertone et al. [7] exhibit a large scatter in the efficiency of this mechanism — the coefficient of variation of the kinetic energy at take off is about $25\%$ for jumps of the same individual.
Given the role of release rate in the biomimetic example of a curved strip on a flat substrate, it is natural to examine the role of release rate in externally latched jumping as a means of controlling the jump characteristics even once the latch has been set. The cleanness of the mechanism of jumping in these examples allows us to build a simple mathematical model of latch-spring-actuated movement that includes normal substrate-system interactions, but neglects friction.
We consider a two-dimensional model of a naturally curved, linearly elastic and inextensible cylindrical shell that lies on the $xz$-plane and both ends are at $z=0$ (see the schematic of the shell in navy blue in the lower inset in fig. 1b). The shell is in contact with a flat, rigid substrate perpendicular to the gravitational field (with acceleration $g$). We shall assume that contact only occurs at these two ends, simplifying the analysis considerably. The natural curvature of the shell is $\kappa_{0}$, and it has width $b$, volumetric mass density $\rho$, Young’s modulus $E$, thickness $h$ and length $\ell$, with $h\ll\ell$.
Initially, an indenting force per unit width $f(t=0)=f_{0}$ is applied at the center of the shell that confines it to lie almost parallel to a planar substrate (the green curve in the schematic of the shell in fig. 1b). This force serves as the motor (increasing and storing elastic energy within the deformed shell for $t<0$). However, once the indenting force reaches $z=0$, $f$ plays the role of the latch, keeping the deformed shell confined — the shell is in equilibrium due to the reaction force of the substrate on the edges. For time $t>0$ the indenting force is gradually released (the teal blue curve in the schematic of the shell in fig. 1b). We assume a simple model for latch dynamics with characteristic unlatching time $t_{u}$,
$f(t)=f_{0}\exp(-t/t_{u}).$ (1)
In the following we non-dimensionalize lengths by $\ell$, heights by $\ell^{2}|\kappa_{0}|$, masses by $\rho h\ell^{2}$, times by $(\rho h\ell^{4}/EI)^{1/2}$ (here $I=h^{3}/12$ is the moment of inertia of the shell per unit width), horizontal forces (per unit width) by $EI/\ell^{2}$ and vertical forces (per unit width) by $EI|\kappa_{0}|/\ell$. We use upper case letters to denote dimensionless versions of the corresponding dimensional quantity, e.g. $X=x/\ell$, $Z=z/(|\kappa_{0}|\ell^{2})$. In the limit of small, planar deflections of a slender shell, the dimensionless vertical coordinates of the shell, $Z(X,T)$, a time $T$ after the gradual release of the force is initiated and before launching are given by the dimensionless dynamic beam equation [8];
$\frac{\partial^{4}Z}{\partial X^{4}}+\frac{\partial^{2}Z}{\partial T^{2}}+H\frac{\partial Z}{\partial T}+1/\beta=0.$ (2)
Here we neglect the tension within the shell, and hence the role of friction with the substrate; this is justified within the small angle approximation where the work done against friction is negligible because of the small horizontal distance moved by the contact points [9].
Two dimensionless groups defined by
$\beta=\frac{EI|\kappa_{0}|}{\rho gh\ell^{2}}\quad\textrm{and}\quad H=\frac{2\eta\ell^{2}}{(EI\rho h)^{1/2}},$ (3)
are introduced. These measure the relative importance of bending forces to the gravitational force and the strength of damping relative to bending forces, respectively. Here $\eta$ is the translational drag coefficient of the shell, and we use slender-body theory to model viscous forces on the shell [10]. In the experiments of fig. 1c, the typical value of $\beta$ is approximately 10, while we estimate it for the larvae of fig. 1a to be $\beta\approx 1$; we must therefore account for gravity in what follows. Similarly, the limit of no damping, $H=0$, is well defined but the decay of oscillations in our experiments suggests that $H\approx 2$.
The applied force $F(T)$ enters via a point loading condition at $X=1/2$ — we assume that there is a discontinuity in the shear force there (or, equivalently, a Dirac delta-function external pressure there). We therefore have that
$\left[\frac{\partial^{3}Z}{\partial X^{3}}\right]_{X=1/2^{-}}^{X=1/2^{+}}=F_{0}\exp(-T/\tau),$ (4)
where $\tau$ is the scaled $t_{u}$ and $F_{0}$ is the dimensionless $f_{0}$ (scaled by the characteristic vertical force per unit width).
The small-slope approximation simplifies the analytic treatment of the problem because, while the shell remains in contact with the substrate, the boundary conditions are fixed at $X=0,1$: $Z(0,T)=Z(1,T)=0$ and $\left.\partial^{2}Z/\partial X^{2}\right|_{X=0,1}=\kappa_{0}/|\kappa_{0}|$. It also avoids energy being dissipated by friction with the substrate (since the edges do not move laterally, there is no work done).
We take as the initial condition that the indentation force is precisely that required to make the center touch the substrate, i.e. $Z(X,T=0)=Z_{0}(X)$ where $Z_{0}(X)$ is the solution to the linearized equilibrium problem
$\frac{{\rm d}^{4}Z_{0}}{{\rm d}X^{4}}+\frac{1}{\beta}=0,$ (5)
with analogous boundary conditions at $X=0,1$ and fixed height at $X=1/2$, $Z_{0}(X=1/2)=0$. (The combination of these conditions yield a unique solution to eq. (5) in the form of a fourth order polynomial that is given in the supplementary information.) The value of $F_{0}$ in (4) is determined from the resulting discontinuity in the shear force. Finally, the shell is assumed to start from rest, so that $\partial Z/\partial T(X,T=0)=0$.
The dynamic beam equation (2) with the boundary, initial and continuity conditions discussed above can be solved analytically by letting
$Z(X,T)=\zeta(X,T)-w(X)-\exp(-T/\tau)\,y(X),$ (6)
where $\zeta(X,T)$ is the separable solution of the homogeneous problem (which depends on the frequencies of natural oscillations $\omega_{n}=\frac{H}{2}\left(\frac{4n^{4}\pi^{4}}{H^{2}}-1\right)^{1/2}$ in underdamped conditions $H<2n^{2}\pi^{2}$), $w(X)$ is a particular solution that absorbs the inhomogeneous parts arising from the natural curvature and $y(X)$ is a particular solution that absorbs the discontinuity imposed by (4). Details of the functional forms for $\zeta(X,T)$, $w(X)$ and $y(X)$ are given in the supplementary information.
Both the jumping of the larvae and our tabletop experiments reveal a broad range of jumping behavior from failed attempts to leave the ground to effective (and impressive) jumps; there is clearly a large variability in the efficiency of this mechanism. We rationalize this behavior in terms of the three dimensionless parameters in the model $\beta,\tau$ and $H$. We begin by noting that the scaled reaction force of the ground on the edge of the shell, $R$, can only be positive — if this reaction force becomes negative then the shell has lost contact with the ground, and the shell has jumped. Noting that $R$ may be calculated from the solution $Z(X,T)$, eq. (6) as
$R(T)=\left.\frac{\partial^{3}Z}{\partial X^{3}}\right|_{X=0}.$ (7)
We then build the phase space of the failure/success in jumping from the analytical solution of the linearized model, eq. (6), finding the smallest positive root of $R(T)=0$ numerically using the Chebfun package [11] in Matlab. The results in fig. 2a show that the shell jumps for larger $\beta$ (light, stiff shell), smaller $\tau$ (quick release) and $H$ (small dissipation), but also that it fails to jump for smaller $\beta$ (heavy shell), larger $\tau$ (slow release) and $H$ (large dissipation).
Figure 2: (a) Phase diagram of the jumping behavior of a curved shell on a flat surface at different values of the damping $H$. The jumping behavior of the shell is characterized by the triad $(\beta,\tau,H)$: there is a jumping phase (in green) and a non-jumping phase (in orange). (b) Phase boundaries at different values of parameter $H$ and small $(\beta,\tau)$. The colour scheme of the boundaries in (a) is the same as in (b).
In undamped conditions ($H=0$), for light, stiff shells (i.e. away from the small $\beta$ limit), the phase boundary in the $(\beta,\tau)$ plane is linear. An asymptotic approximation of the analytical solution eq. (6) — obtained by truncating at the fundamental mode $(n=1)$ and considering the limit of large $\tau$ and large $T/\tau$ — allows us to calculate an approximation of $R(T)$ and, upon imposing the condition that the minima in this approximation equal to zero, we find an estimate of the phase boundary $\tau_{\text{boundary}}(\beta)\sim\frac{24}{\pi^{3}}\beta$ (a detailed derivation is provided in the supplementary information) that shows good agreement with numerical results for $\beta>1$ (see fig. 2b). This boundary becomes noticeably sublinear as the dissipation $H$ increases: jumping requires a quicker unlatching for more strongly damped systems. Moreover, regardless of the damping $H$, a strong nonlinear deviation of the phase boundary also occurs for small $\beta$ (fig. 2b). The phase boundary crosses the abscissa at $\beta\sim\beta_{\text{crit}}$, where $\beta_{\text{crit}}=\frac{1}{2}\frac{\left(\frac{5}{2}+\pi\right)}{12+\pi}$ is an estimation (by asymptotic arguments, see the supplementary information) of the minimal $\beta$ for jumping. This minimum is motivated by the fact that a sufficiently heavy shell will bend due to its own weight and hence naturally satisfy the condition $Z_{0}(X=1/2)=0$ at zero external force.
The phase space in fig. 2 quantifies the intuition that a faster unlatching mechanism facilitates jumping. We therefore expect that a jumping mechanism should be designed to ensure the fastest possible unlatching. However, the larvae studied by Bertone et al. [7] seem to contradict this basic principle as they grip the substrate with all their legs, which presumably results in a slower latch release. To understand this apparent paradox we examine the maximum dimensional latch force $F_{0}$ — i.e. the attaching force between the larva’s legs and the substrate — and the stored energy (bending + gravitational). The results of this calculation show (fig. 3) that the magnitude of the latch force increases approximately linearly as the curvature increases — even when geometrical nonlinearities are accounted for. However, the stored energy shows a superlinear increase with curvature (see SI fig. S6). Detailed experiments on other insect larvae [12] show that the legs are typically able to supply a maximum adhesive force of $0.3-0.6\mathrm{\penalty 10000\ mN}$. Assuming that each pair of legs is able to produce a similar force, and noting that they are located nearby one another (and hence can be assumed to be approximately a point force), using multiple legs to anchor on the surface should increase the maximum natural curvature obtained before release by a factor of 3 and the energy stored by a factor of 9. We therefore estimate the maximum latch force that can be generated by Laemophloeus biguttatus should be around $0.2\mathrm{\penalty 10000\ mN}$ (approximating its body by a rod of radius $\sim 0.6\mathrm{\penalty 10000\ mm}$). In our model, this corresponds to a maximum natural curvature $\kappa_{0}^{\max}$, but to determine the corresponding dimensional value requires an estimate of the effective Young’s modulus of the larva. We obtain such an estimate by using the time scale of the return to the naturally curved state once the jump has been initiated, and comparing this to theoretical work on the curling of a ribbon by Callan-Jones et al. [13]; this leads us to estimate $E_{\rm eff}\sim 10\mathrm{\penalty 10000\ kPa}$, and corresponds to $|\kappa_{0}^{\max}|=1.69\mathrm{\penalty 10000\ mm}^{-1}$ (see SI). Figure 3 shows that these estimates are consistent. Finally, we hypothesize that once the maximum force is exceeded at one point and the adhesion between leg and substrate fails, then the remaining legs are successively detached.
Figure 3: Latch force required for an indented, naturally curved, shell with dimensions from Laemophloeus biguttatus [7]. Latch force as a function of the shell’s curvature given by the linearized theory (in red) and the non-linear model (in blue). Line styles denote different Young’s modulus: $E\sim 10\penalty 10000\ \mathrm{kPa}$ (estimated value from the time scale of curling of an elastic object [13], solid lines); $E\sim 9.25\penalty 10000\ \mathrm{kPa}$ (estimated value from the attachment force of 2nd instar of G. viridula [12], dashed lines); $E\sim 14.9\penalty 10000\ \mathrm{kPa}$ (estimated value from the attachment force of 3rd instar of G. viridula [12], dot–dashed lines). (Here ‘instar’ refers to different developmental stages of insects, between moulting of the exoskeleton.)
A key parameter of importance in the space of LaMSA mechanisms is the efficiency of the release mechanism: what percentage of the elastic energy that is initially stored is realized as translational kinetic energy upon release? Yang and Kim [5] measured this for elastic jumping hoops and found that it is a universal quantity, around $57\%$, independent of the material properties of their system. While it is difficult to ascribe a precise value of $\tau$ to their experimental technique, it is very small, and agrees well with the predicted efficiency when $\tau\approx 0$. Here, we are able to readily calculate the efficiency of jumping, defined as
$\mathcal{E}=\frac{K_{\rm trans}}{E_{0}}\times 100\%$ (8)
directly from the solution in (6). Our results for $\mathcal{E}(\tau)$ are shown for a variety of values of $\beta$ and $H$ in fig. 4. As might have been expected, these show that the efficiency of jumping decreases with $\tau$, though it is somewhat surprising that our results suggest that this decay appears to be exponential. Moreover, there appear to be certain pessimal values of $\tau$ where the efficiency drops even more dramatically (we estimate these pessimal values in the SI using asymptotic methods). We attribute this to a destructive interference between the imposed decay of $F(t)$ and the natural relaxation time of the shell, which is consistent with the re-entrant efficiency observed for lighter, stiffer shells.
The sensitive dependence of the efficiency of jumping on the release time scale $\tau$ suggests a rationalization for the wide variability in the efficiency of jumps that were reported by Bertone et al. [7]: even an apparently small change in the latch release time $\tau$ can lead to a large decrease in the efficiency of jumping. This complements previous studies, which have focussed on the role of latch geometry and release velocity in control of energy release with internal latches [14, 15].
Figure 4: Energy efficiency of jumping of a curved shell given by the ratio of translational kinetic energy at $T=T_{\text{off}}$ to stored energy at $T=0$ as a function of the latch lifting time scale. Theoretical predictions from linearized theory are depicted in solid lines. Scaled values of the efficiency for L. biguttatus [7], jumping hoops [5], and our experiments of a curved strip are also included as symbols. Solid lines for $\beta=1$ and dashed lines for $\beta=10$. Colors range from dark brown (small $H$) to yellow (large $H$) for $H=0.1,1,2,10$,
We have developed a dynamical theory that describes the observed efficiency of the jumps of insect larvae, as well as providing a mechanism for the wide distribution of observed jump efficiencies: an exponential decay means that small changes in $\tau$ lead to significantly less efficient jumps. While this may pose a substantial challenge to insect larvae, it also focuses the design criteria for controlling jumps in robots [6], for example. To make this a reality, however, may require generalizing this work to scenarios where the contact set is more complicated than that considered here — for example, in insect-inspired jumping robots, the contact region grows during the early stages of jumping, before ultimately decreasing [6].
Acknowledgements.
This work was partially supported by the UK Engineering and Physical Sciences Research Council via Grant No. EP/W016249/1 and Grant No. EP/Y027949/1 (M.S. and D.V.), and by the Swedish Research Council under Grant No. 638-2013-9243 (M.S. and J.W.). For the purpose of Open Access, the authors will apply a CC BY public copyright license to any Author Accepted Manuscript version arising from this submission.